OpenStudy (anonymous):

find the domain algebraically, f(x)=2x^3-5x+6

myininaya (myininaya):

This is a polynomial the domain is all read numbers

OpenStudy (anonymous):

\[\left\{ x | x \in \mathbb{R} \right\}\]

myininaya (myininaya):

i like \[(-\infty,\infty)\] better ;) so i win

OpenStudy (anonymous):

I need to see the work on this